OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the y's to add out? 3x - y = 3 -2x + 2y = 6 a)3 b)2 c)6 d)none of the above

OpenStudy (mathstudent55):

Look at the second equation. In it, y has a coefficient of 2. The y term of the second equation is 2y. What do you add to 2y to get zero? That's what you need to have in the first equation. Since the first equation has a y term of -y, you need to multiply it by a number to make it such that when you add it to 2y you get zero. Since you need to multipy the entire equation by that number, that's why the question asks what number you need to multiply the first equation by, but the way to tell is to figure out what you need to multiply y of the first equation by.

OpenStudy (anonymous):

so the answers is 2? I got 6

OpenStudy (mathstudent55):

Yes, the answer is 2. You need to multiply the first equation buy 2 to get a y term of 2y, since 2y plus the -2y of the second equation will add up to zero. This is not the solution of the system of equations. All they are asking is what to multiply the first equation by.

OpenStudy (anonymous):

ooooh, can you walk me through another problem just like this one?

OpenStudy (mathstudent55):

This is the explanation of how the addition method works. 3x - y = 3 multiply equation by 2 ------> 6x - 2y = 6 -2x + 2y = 6 keep the same ------------> -2x + 2y = 6 Now add the equations -------------> 4x = 12 x = 3 Now substitute 3 in for x in first original equation: 3(3) - y = 3 9 - y = 3 -y = -6 y = 6 The solution is x = 3, y = 6 Keep in mind, this is not what you were asked for. All you were asked for was what number do you multiply the first equation by to eliminate the y term using the addition method. That answer is 2.

OpenStudy (anonymous):

To solve this system using the addition method, you would need to multiply the first equation by what number in order for the x's to add out? 2x + 3y = 8 -6x + 4y = 11

OpenStudy (anonymous):

for this probblem i tired following your step but i got confused

OpenStudy (mathstudent55):

This is similar to the prevous problem. What do you add to -6x to get zero?

OpenStudy (anonymous):

you'll have to divide ?

OpenStudy (mathstudent55):

Answer my question above.

OpenStudy (mathstudent55):

-6x + ______ = 0 what goes in the blank?

OpenStudy (anonymous):

sorry i was on the phone real quick, but yea 6

OpenStudy (mathstudent55):

No, you mean 6x, right?

OpenStudy (anonymous):

@mathstudent55 its 6

OpenStudy (mathstudent55):

No again. 6 + (-6x) is not equal to zero. What do you add to -6x to get zero?